The car starts moving along the direction with constant acceleration. In the first second of his movement, he crosses 2m. How much extra time will he cover in the second second of his movement?


1
Expert's answer
2021-03-21T11:24:16-0400

Let's first find the acceleration of the car:


"d=v_0t+\dfrac{1}{2}at^2."

Since the car starts from rest, we get:


"d=\dfrac{1}{2}at^2,""a=\dfrac{2d}{t^2}=\dfrac{2\cdot2\ m}{(1\ s)^2}=4\ \dfrac{m}{s^2}."

Finally, we can find the distance covered by the car in the 2 s of his movement:


"d(t=2\ s)=\dfrac{1}{2}\cdot4\ \dfrac{m}{s^2}\cdot(2\ s)^2=8\ m."